OpenStudy (anonymous):

A sum of money amounting to $5.80 consists of dimes and quarters. If there are 34 coins in all, How many are quarters ?

OpenStudy (aaronq):

let x be the number of quarters and y the number of dimes 34 coins in all: x + y = 34 amounting to $5.80: 0.25x+0.1y=5.80 solve for x, the number of quarters

OpenStudy (anonymous):

im confused on how do i solve for x

OpenStudy (anonymous):

is it 22 ?

OpenStudy (anonymous):

21 rounded

OpenStudy (aaronq):

isolate y, x + y = 34 -> y=34-x substitute into 0.25x+0.1y=5.80 0.25x+0.1(34-x)=5.80 x=16

OpenStudy (anonymous):

thanks

OpenStudy (aaronq):

no problem

OpenStudy (anonymous):

i just put another one up same type of problem